OpenStudy (anonymous):

I f x is an integer which of the following must be an even integer? a. x^2-x-1 b. x^2-4x+6 c. x^2-5x+5 d. x^2+3x+8 e. x^2+2x+10

OpenStudy (zarkon):

d

OpenStudy (anonymous):

why though?

OpenStudy (zarkon):

odd * odd = odd even * even = even even +even=even odd+odd =even

OpenStudy (zarkon):

for x^2+3x+8 is x is odd then x^2 is odd and 3x is odd odd+odd+even=even if x is even then x^2 is even and 3x is even even+even+even=even

OpenStudy (zarkon):

you could also use x=2k+1 and 2k and see that both are even in (d)

myininaya (myininaya):

if x is even, then x=2n for some integer n so x^2-x-1=(2n)^2-2n-1=(2n)^2+(2n+1)=2(2n^2)+(2n+1)=even+odd=even but what if x is odd, then x=2n+1 for some integer n so x^2-x-1=(2n+1)^2-(2n+1)-1=(2n+1)(2n+1-1)-1=(2n+1)(2n)-1=2n(2n+1)-1 =even-1=odd

myininaya (myininaya):

so choice a it like totally not right since we got an odd output

OpenStudy (zarkon):

let x=0 to see (a) does not work

myininaya (myininaya):

or x=3 9-3-1=6-1=5

myininaya (myininaya):

but 0 is pretty

myininaya (myininaya):

prettier*

OpenStudy (anonymous):

thanks guys